OpenStudy (anonymous):

This is the word problem I have to solve. Tyler is twice as old as his son. ten years ago Tyler was three times as old as his son. How old are they? Must show work

OpenStudy (anonymous):

Let tyler's age be x, and his son's, y. Then relate different equations as a function of time.

OpenStudy (anonymous):

let t = tyler's age let s = son's age t = 2s t - 10 = 3(s - 10) 2s - 10 = 3s - 3- s = 20 son is 20 tyler is 40

OpenStudy (anonymous):

manderson do you use only 1 variable x or are you studying how to use 2 variables x and y?